find the value of m such that this question have no solutions:
3x-my=4
x+y=12

Answer by ichudov(507)

that's a great problem. Start solving this system a little bit, it will help. Subtract 3 times the bottom equation, from top equation:
3x-my - 3(x+y) = 4-3*12
-my-3y = -32
remove minuses just to make it more readable
my+3y = 32
the value of m that makes this equation impossible (no solutions) is such am that makes the equation 0=32. 0=32 is an impossible equation.
-3 looks pretty good:
-3y+3y=32 is the same as 0 = 32.
so, the answer is m=-3.